OpenStudy (anonymous):

show solution please... @lgbasallote , \[\large \int_{0}^{s}axdx= \]

OpenStudy (anonymous):

a is a constant...

OpenStudy (anonymous):

as^2 /2 ?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[\large \int_{0}^{s}axdx=[\frac{ax^2}{2}]^{s} _{0} =\frac{as^2}{2}\]

OpenStudy (unklerhaukus):

\[\int\limits_0^s ax\cdot\text dx=\left.\frac{ax^2}{2}\right|_0^s=\frac{as^2}{2}\]

OpenStudy (anonymous):

ok... that makes sense...

OpenStudy (unklerhaukus):

yes?

OpenStudy (anonymous):

perfect.. thank you...:)

OpenStudy (unklerhaukus):

too easy

OpenStudy (anonymous):

Not easy for me..

OpenStudy (anonymous):

me neither.... someone told me this was a special integral... can't remember though...

OpenStudy (anonymous):

a is a constant => take it out of the integral => \(\large a \int_{0}^{s}xdx\) Then, it looks good :)

OpenStudy (anonymous):

so you can pull any constant out of the integral sign , right?

OpenStudy (unklerhaukus):

yes you can factor a constant out of the integral

OpenStudy (anonymous):

Hmm.. Sorry, I should make it clear.. The coefficient of variable. Constant is like +C => not pull out..

OpenStudy (anonymous):

so why was this supposed to be some special integral? idk....

OpenStudy (anonymous):

Perhaps, it's special when you have to apply it? I don't know too, sorry!

OpenStudy (unklerhaukus):

\[\int\limits_a^b (x+a)\cdot\text dx=\int\limits_a^b x\cdot \text dx+a\int\limits_a^b \text dx=\left.\left(\frac{x^2}{2}+ax\right)\right|_a^b\]

OpenStudy (anonymous):

@lgbasallote ,??

OpenStudy (anonymous):

Oh.. that a is a constant!!

OpenStudy (saifoo.khan):

@dpaInc asked a question! :O Am i dreaming?

OpenStudy (anonymous):

@saifoo.khan You're not

OpenStudy (saifoo.khan):

Pinch me!

OpenStudy (lgbasallote):

\[\huge \frac{as^2}{2} = \frac{a*s*s}{2}\]

OpenStudy (anonymous):

:O, half-a** integral...

OpenStudy (lgbasallote):

another fun integral \[\huge \int e^{1+ \ln x}dx\] try it :)

OpenStudy (anonymous):

tough one...

OpenStudy (anonymous):

\[\int e^{1+ \ln x}dx = \int e(e^{\ln x})dx = \int exdx = e\int xdx = ...\]

OpenStudy (anonymous):

why can you pull the e out?

OpenStudy (anonymous):

because e is a constant, and its value is 2.718